# Beyond Seen Data: Improving KBQA Generalization Through Schema-Guided Logical Form Generation

### v2

@@ 5.3 Ablation Study (Q2)
Model GrailQA WebQSP
Overall I.I.D. Comp. Zero. Overall
SG-KBQA 88.5 94.6 84.6 87.9 80.3
w/o RG-EMD 85.3 92.4 80.2 84.3 78.4
w/o RG-CER 86.5 92.1 81.1 86.3 79.5
w/o DED 87.8 94.0 82.4 87.2 78.2
w/o SC 79.2 92.9 77.4 73.9 77.1
Table 3: Ablation study results (F1 score) on the validation set of GrailQA and the test set of WebQSP.
Next, we run an ablation study with the following variants of SG-KBQA: w/o RG-EMD replaces our relation-guided entity mention detection with SpanMD Shu et al. (2022) which is commonly used in existing models Pang et al. (2022); Faldu et al. (2024); w/o RG-CER omits candidate entities retrieved from the top relations; w/o DED uses the top-1 candidate entity for each entity mention without deferring entity disambiguation; w/o SC omits schema contexts from logical form generation.
Table 3 shows the results on the validation set of GrailQA and the test set of WebQSP. Only F1 scores are reported for conciseness, as the EM scores on GrailQA exhibit similar comparative trends and are provided in Appendix F.
All model variants have lower F1 scores than those of the full model, confirming the effectiveness of the model components. SG-KBQA w/o DED (with schema contexts) reduces the F1 scores across various generalization settings on both datasets, demonstrating the effectiveness of our DED strategy in reducing error propagation during the retrieval and generation stages. Furthermore, SG-KBQA w/o SC (with deferred entity disambiguation) has the most significant drops in the F1 score under the compositional (7.2) and zero-shot (14.0) generalization tests. It highlights the importance of schema contexts in constraining the larger search space introduced by DED and in generalizing to unseen KB elements and their combinations. Meanwhile, the lower F1 of SG-KBQA w/o RG-EMD emphasizes the capability of our relation-guided entity mention detection module in strengthening KBQA entity retrieval.
@@ 5.4 Module Applicability (Q3)
Our relation-guided entity retrieval (RG-EMD & RG-CER) module and schema-guided logical form generation (DED & SC) module can be applied to existing KBQA models. We showcase such applicability with the TIARA model. As shown in Table 4, by replacing the retrieval and generation modules of TIARA with ours, the F1 scores increase consistently for the non-I.I.D. tests.
Table 4 further reports F1 scores of SG-KBQA when we replace LLaMA3.1-8B with T5-base (which is used by TIARA), and DeepSeek-R1-Distill-Llama-8B (DS-R1-8BGuo et al. (2025) for logical form generation. We see that, even with the same T5-base model for the logical form generator, SG-KBQA outperforms TIARA consistently. This further confirms the effectiveness of our model design. As for DS-R1-8B, it offers accuracy slightly lower than that of the default LLaMA3.1-8B model. We conjecture that this is because DS-R1-8B is distilled from DeepSeek-R1-Zero, which focuses on reasoning capabilities and is not specifically optimized for the generation task.
We also have results on parameter impact, model running time, a case study, and error analyses. They are documented in Appendices H to K.
Model Overall I.I.D. Comp. Zero.
TIARA (T5-base) 81.9 91.2 74.8 80.7
w RG-EMD & RG-CER 84.3 92.3 78.1 83.3
w DED & SC 85.6 92.3 79.8 85.0
SG-KBQA 88.5 94.6 83.6 87.9
w T5-base 84.9 92.6 81.0 83.3
w DS-R1-8B 87.5 94.0 82.4 86.7
Table 4: Module applicability results (F1 score) on the validation set of GrailQA. EM scores are in Appendix G.
@@ 6 Conclusion
We proposed SG-KBQA for the KBQA task. Our core innovations include: (1) using relation to guide the retrieval of entities; (2) deferring entity disambiguation to the logical form generation stage; and (3) enriching logical form generation with schema contexts to constrain search space. Together, we achieve a model that tops the leaderboard of a popular non-I.I.D. dataset GrailQA, outperforming SOTA models by 4.0%, 2.9%, and 3.3% in F1 under compositional generalization, zero-shot generalization, and overall test settings, respectively. Our model also performs well in the I.I.D. setting, outperforming SOTA models on WebQSP.
@@
First, like any other supervised models, SG-KBQA requires annotated samples for training which may be difficult to obtain for many domains. Exploiting LLMs to generate synthetic training data is a promising direction to address this issue. Second, as discussed in the error analysis in Appendix K, errors can still arise from the relation retrieval, entity retrieval, and logical form generation modules. There are rich opportunities in further strengthening these modules. Particularly, as we start from relation extraction, the overall model accuracy relies on highly accurate relation extraction. It would be interesting to explore how well SG-KBQA performs on even larger KBs with more relations.
@@
Appendix A S-Expression
@@
a set of entities These functions return all $x$ such that $(x,v)\in b$ and $v$ $<$ / $\leq$ / $>$ / $\geq$ $n$
@@
Figure 5: Example prompt to our fine-tuned LLM-based logical form generator for an input question: Captain pugwash makes an appearance in which comic strip?
@@ Appendix B Prompt Example
We show an example prompt to our fine-tuned LLM-based logical form generator containing top-20 relations and top-2 entities per mention retrieved by our model in Figure 5.
Appendix C Additional Details on the WebQSP Dataset
WebQuestionsSP (WebQSP) Yih et al. (2016) is an I.I.D. dataset. It contains 4,937 questions collected from Google query logs, including 3,098 questions for training and 1,639 for testing, each annotated with a target SPARQL query. We follow GMT-KBQA Hu et al. (2022b), TIARA Shu et al. (2022) to separate 200 questions from the training questions to form the validation set.
Appendix D Baseline Models
@@
  • RetinaQA Faldu et al. (2024) uses both a ranking-based method and a generation-based method (TIARA) to generate logical forms, which are then scored by a discriminative model to determine the output logical form.
@@
  • ChatKBQA Luo et al. (2024) fine-tunes an open-source LLM to map questions into draft logical forms. The ambiguous KB items in the draft logical forms are replaced with specific KB elements by a separate retrieval module.
@@
Overall I.I.D. Compositional Zero-shot
Model EM F1 EM F1 EM F1 EM F1
SG-KBQA 85.1 88.5 93.1 94.6 78.4 83.6 84.4 87.9
[…]
   w/o RG-CER 82.8 86.5 90.2 92.1 75.4 81.1 82.7 86.3
   w/o DED 84.3 87.8 92.6 94.0 77.1 82.4 83.7 87.2
   w/o SC 76.6 79.2 91.7 92.9 72.3 77.4 71.7 73.9
   w/o Fallback LF 81.8 84.6 92.8 94.1 77.3 81.8 78.7 81.5
Table 6: Ablation study results on the validation set of GrailQA.
Appendix E Implementation Details
@@
Appendix F Full Ablation Study Results (GrailQA)
Table 6 presents the full ablation study results on the validation set of GrailQA. We observe a similar trend to that of the F1 score results reported earlier – all ablated model variants yield lower EM scores compared to the full model.
For the retrieval modules, RG-EMD improves the F1 score by 3.2 points and the EM score by 3.8 points on GrailQA (i.e., SG-KBQA vs. SG-KBQA w/o RG-EMD for overall results), while achieving a 1.9-point increase in the F1 score on WebQSP (see Table 2 earlier). It achieves an increase of 3.4 points or larger in the F1 score on the compositional and zero-shot tests, which is larger than the 2.2-point improvement on the I.I.D. tests. This shows that relation-guided mention detection effectively enhances the generalization capability of KBQA entity retrieval. For the other module RG-CER, removing it (SG-KBQA w/o RG-CER) results in a 2.5-point drop in the F1 score for both the I.I.D. and compositional tests, while the impact is smaller on the zero-shot tests (1.6 points). This is because the lower accuracy in relation retrieval under zero-shot tests leads to error propagation into relation-guided candidate entity retrieval, reducing the benefits of this module.
For the generation modules, SG-KBQA w/o DED negatively impacts the F1 scores on both GrailQA and WebQSP, confirming that deferring entity disambiguation effectively mitigate error propagation between the retrieval and generation stages. For SG-KBQA w/o SC, it reduces the F1 score by 1.7 points and 3.2 points on the GrialQA I.I.D. tests and on WebQSP. The drop is more significant on the compositional and zero-shot tests, i.e., by 6.2 points and 14.0 points, respectively. This indicates that schema contexts can effectively guide the LLM to reason and identify the correct combinations of KB elements unseen at training.
In Table 6, we present an additional model variant, SG-KBQA w/o Fallback LF, which removes the fall back logical form generation strategy from SG-KBQA. We see that SG-KBQA has lower accuracy without the strategy. We note that this fallback strategy is not the reason why SG-KBQA outperforms the baseline models. TIARA also uses this fallback strategy, while RetinaQA uses the top executable logical form from the fallback strategy as one of the options to be selected by its discriminator to determine the final logical form output.
Appendix G Full Module Applicability Results
Overall I.I.D. Compositional Zero-shot
Model EM F1 EM F1 EM F1 EM F1
TIARA (T5-base) 75.3 81.9 88.4 91.2 66.4 74.8 73.3 80.7
   w RG-EMD & RG-CER 79.5 84.3 90.3 92.3 71.2 78.1 78.3 83.3
   w DED & SC 79.9 85.6 88.6 92.3 72.7 79.8 79.0 85.0
SG-KBQA 85.1 88.5 93.1 94.6 78.4 83.6 84.4 87.9
   w T5-base 80.6 84.9 89.9 92.6 73.8 81.0 79.4 83.3
   w DS-R1-8B 83.6 87.5 92.3 94.0 75.4 82.4 83.1 86.7
Table 7: Full module applicability results on the validation set of GrailQA.
To evaluate the applicability of our proposed modules, we conduct a module applicability study with TIARA (an open-source retrieve-then-generate baseline) and different generation models (i.e., T5-base and DeepSeek-R1-Distill-Llama-8B).
Table 7 reports the results. Replacing TIARA’s entity retrieval module with ours (TIARA w RG-EMD & RG-CER) helps boost the EM and F1 scores by 4.2 and 2.4 points overall, comparing against the original TIARA model. This improvement is primarily from the tests with KB elements or compositions that are unseen at training, as evidenced by the larger performance gains on the compositional and zero-shot tests, i.e., 3.3 and 2.6 points in the F1 score, respectively. Similar patterns are observed for TIARA w DED & SC that replaces TIARA’s logical form generation module with ours. These results demonstrate that our proposed modules can enhance the retrieval and generation steps of other compatible models, especially under non-I.I.D. settings.
Further, using the same language model (i.e., T5-base in TIARA) to form logical form generation modules, our model SG-KBQA w T5-base still outperforms TIARA by 5.3 points 3.0 points in the EM and F1 scores for the overall tests. This confirms that the overall effectiveness of our model stems from its design rather than the use of a larger model for logical form generation. As for SG-KBQA w/ DS-R1-8B, it reports close performance to SG-KBQA, indicating that SG-KBQA does not rely on a particular LLM.
Appendix H Parameter Study
We conduct a parameter study to investigate the impact of the choice of values for our system parameters. When the value of a parameter is varied, default values as mentioned in Appendix E are used for the other parameters.
@@ Appendix H Parameter Study
As $k_{L}$ increases, the recall of candidate entity retrieval grows, which is expected. The growth diminishes gradually. This is because a small number of questions contain complex entity mentions that are difficult to handle (see error analysis in Appendix K). As $k_{L}$ increases, the precision of the retrieval also reduces, which brings noise into the entity retrieval results and additional computational costs. To strike a balance, we set $k_{L}=3$ for GrailQA and $k_{L}=4$ for WebQSP. We also observe that the recall on WebQSP is lower than that on GrailQA. This is because WebQSP has a smaller training set to learn from.
@@
SpanMD: What is the name for the atomic units of length? (✗)
Ours:
@@
Appendix I Model Running Time
@@
Appendix J Case Study
Question: Captain pugwash makes an appearance in which comic strip?
Relation Retrieval Entity Retrieval
TIARA …comic_strips_appeared_in Captain Pugwash m.04fgkzf
…character
(AND comic_strips.comic_strip_character (JOIN
comic_strips.comic_strip_character.comic_strips_appeared_in m.04fgkzf))
Ours [D] comic_strip_character [ID] m.04fgkzf
[N] comic_strips_appeared_in [N] Captain Pugwash
[R] comic_strip [C] comic_strip
@@
[N] character [N] Captain Pugwash
[R] comic_strip_character [C] comic_strip_character
(AND comic_strips.comic_strip (JOIN comic_strips.comic_strip.characters m.02hcty))    (✓)
Table 9: Case study of logical form generation by SG-KBQA and a representative competitor TIARA on the GrailQA validation set. Incorrect relations and entities are marked in red, while the correct relations and entities are colored in blue.
To further show SG-KBQA’s generalizability to non-I.I.D. KBQA applications, we include a case study from the GrailQA validation set as shown in Tables 8 and 9.
Entity Mention Detection
Figure 8 shows an entity mention detection example, comparing our entity detection module with SpanMD which is a mention detection method commonly used by SOTA KBQA models Shu et al. (2022); Ye et al. (2022); Faldu et al. (2024). In this case, SpanMD incorrectly detects length as an entity mention, which is actually part of the ground-truth relation (measurement_unit.$\ldots$.length_units) that is unseen in the training data. Our entity mention detection module, on the other hand, leverages the retrieved relations to generate a logical form sketch. The correct entity mention, atomic units, is isolated from the relations and can be corrected extracted, even though this entity mention has not been seen at training.
@@
Table 9 shows a logical form generation example. Here, SG-KBQA and TIARA (a representative generation-based model) have both retrieved the same sets of relations in the retrieval stage which include false positives. The two models also share the same top-1 retrieved entity m.04fgkzf, while SG-KBQA has retrieved a second entity m.02hcty in addition. TIARA is misled by the erroneous KB relations retrieved and produces an incorrect logical form. SG-KBQA, on the other hand, is able to produce the correct logical form by leveraging the schema information (i.e., the entity’s class and the relation’s domain and range classes).
Appendix K Error Analysis

### v3

@@ 5.3 Ablation Study (Q2)
Next, we run an ablation study with the following variants of SG-KBQA: w/o SER replaces our schema-guided entity retrieval with the entity linking results from TIARA (Shu et al., 2022) which is commonly used in the baselines (Gu et al., 2023; Faldu et al., 2024); w/o DED uses the top-1 candidate entity for each entity mention without deferring entity disambiguation; w/o SC omits schema contexts (classes) from logical form generation; w/o Fallback LF removes the fall back logical form generation mechanism from SG-KBQA. Table 3 shows the results on the validation set of GrailQA and the test set of WebQSP.
Schema-guided Entity Retrieval
SER improves the F1 score by 3.7 points and the EM score by 4.1 points on GrailQA (i.e., SG-KBQA vs. SG-KBQA w/o SER for overall results), while achieving a 1.9-point increase in the F1 score on WebQSP. It yields an improvement of at least 3.9 F1 points and 4.5 EM points on the compositional and zero-shot test sets, which exceeds the gains observed on the I.I.D. test set (2.5 F1 points and 3.0 EM points). This indicates that incorporating retrieved relations helps the model more accurately identify them and entity mention boundaries in questions involving unseen relations. Since the entity retrieval results directly serve as input to the generation stage, this improvement further enhances the model’s overall generalization performance under non-I.I.D. settings. More discussion on entity retrieval results and a case study are provided in Appendix J and Appendix H, respectively.
Schema-guided Logical Form Generation
SG-KBQA w/o DED negatively impacts the F1 scores on both GrailQA and WebQSP, confirming that our DED strategy effectively mitigate error propagation between the retrieval and generation stages. Meanwhile, SG-KBQA w/o SC (with deferred entity disambiguation but no class information) has the most significant drops in F1 under the compositional (7.2) and zero-shot (14.0) tests. This highlights the contribution of class information in enabling the model to understand the connections among retrieved KB elements, thereby facilitating the generation of correct logical forms. A case study illustrating the schema-guided logical form generation module is provided in Appendix H.
Fallback Mechanism
SG-KBQA w/o Fallback LF exhibits slightly lower performance, with F1 decreasing by 0.9 and EM by 3.3 on GrailQA, and F1 decreasing by 1.7 on WebQSP. We note that this fallback mechanism is not the reason why SG-KBQA outperforms the baseline models. TIARA also uses this fallback mechanism, while RetinaQA uses the top executable logical form from the fallback mechanism as one of the options to be selected by its discriminator to determine the final logical form output. As shown in Table 3, SG-KBQA still produces competitive results without fallback on both I.I.D and non-I.I.D settings. Figure 4 further reports the proportion of questions where the fallback mechanism was triggered, which is at most 11.1% of the test questions. This demonstrates that our SG-KBQA model can generate executable logical forms under non-I.I.D. settings in most cases without relying on fallback.
Figure 4: Proportion of questions resolved by the fallback mechanism on the validation set of GrailQA and the test set of WebQSP.
@@ 5.4 Module Applicability (Q3)
Our entity retrieval module SER and logical form generation module DED & SC can be applied to existing models to improve their generalizability under non-I.I.D. settings. We showcase such applicability with TIARA. As shown in Table 4, replacing TIARA’s entity retrieval module with ours (TIARA w SER) helps boost the EM and F1 scores by 4.2 and 2.4 points overall, comparing against the original TIARA model. This improvement is primarily from the tests with KB elements or compositions that are unseen at training, as evidenced by the larger performance gains on the compositional and zero-shot tests, i.e., 3.3 and 2.6 points in F1, respectively. Similar patterns are observed for TIARA w DED & SC that replaces TIARA’s logical form generation module with ours. These results demonstrate that our modules can enhance the retrieval and generation steps of other compatible models, especially under non-I.I.D. settings.
Table 4 further reports EM and F1 scores of SG-KBQA when replacing LLaMA3.1-8B with T5-base (which is used by TIARA) for logical form generation. We see that, even with the same T5-base model for the logical form generator, SG-KBQA outperforms TIARA by 5.3 points 3.0 points in the EM and F1 scores for the overall tests. This further confirms that the performance gains come from the incorporation of the class contexts instead of a more advanced backbone model.
We also have results on parameter impact, model running time, a case study, and error analyses. They are documented in Appendices F to I.
@@ 6 Conclusion
We proposed SG-KBQA for KBQA. Our core innovations include: (1) using relation to guide entity retrieval; (2) deferring entity disambiguation to the logical form generation stage; and (3) enriching logical form generation with schema (class) contexts indicate KB element connections. Together, we achieve a model that tops the leaderboard of a popular non-I.I.D. dataset GrailQA, outperforming SOTA models by 4.0%, 2.9%, and 3.3% in F1 under compositional generalization, zero-shot generalization, and overall test settings, respectively. Our model also performs well in the I.I.D. setting, outperforming SOTA models on WebQSP and CWQ.
@@
Like any other supervised models, SG-KBQA requires annotated samples for training which may be difficult to obtain for many domains. Exploiting LLMs to generate synthetic training data is a promising direction to address this issue.
Also, as discussed in the error analysis in Appendix I, errors can still arise from the relation retrieval, entity retrieval, and logical form generation modules. There are rich opportunities in further strengthening these modules. As we start from relation extraction, the overall model accuracy relies on highly accurate relation extraction. It would be interesting to explore how well SG-KBQA performs on even larger KBs with more relations.
We further noted several recent works in this highly competitive area, e.g., READS (Xu et al., 2025a) and MemQ (Xu et al., 2025b), which fine-tune LLMs for reasoning trajectory generation or step-wise tool invocation based on given topic entities. These studies represent parallel efforts to ours with a focus on I.I.D. settings. We plan to evaluate their performance in future work, particularly under settings where the topic entity is unavailable and in our non-I.I.D. scenarios.
@@
Acknowledgments
This work is in part supported by the Australian Research Council (ARC) via Discovery Project DP240101006. Jianzhong Qi is supported by ARC Future Fellowship FT240100170.
Appendix A Basic Concepts
Figure 5: A subgraph of Freebase (top) and its corresponding ontology (bottom).
Ontology
As shown in Figure 5, an ontology triple example is:
book.literary_series,
book.literary_series.author,
book.author
An instance of it is:
Harry Potter,
book.literary_series.author,
J.K. Rowling
Here, Harry Potter is an entity that belongs to class book.literary_series; J.K. Rowling is an entity that belongs to class book.author.
S-expressions
@@ Appendix A Basic Concepts
a set of entities These functions return all $x$ such that $(x,v)\in b$ and $v$ << / $\leq$ / >> / $\geq$ $n$
@@
Example Prompt
Please translate the following question into logical form using the provided relations and entities.
Question: Strong lyrics is the description of which video game rating?
Candidate relations with their corresponding Domain [D], Name [N], Range [R]:
[D] cvg.game_version
[N] cvg.game_version.rating
[R] cvg.computer_game_evaluation;
[D] cvg.computer_game_rating
[N] cvg.computer_game_rating.rating_system
[R] cvg.computer_game_rating_system;
[D] cvg.computer_game_rating_system
[N] cvg.computer_game_rating_system.content_descriptors
[R] cvg.computer_game_content_descriptors
…(Continue in the same manner for additional relations)
Candidate entities with their corresponding id [ID], Name [N], Class [C]:
[ID] m.042zlv3
[N] Strong lyrics
[C] cvg.computer_game_content_descriptors
…(Continue in the same manner for additional entities)
Table 6: Example prompt to our fine-tuned LLM-based logical form generator for an input question: Strong lyrics is the description of which video game rating?
@@ Appendix B Prompt Example
We show an example prompt to our fine-tuned LLM-based logical form generator containing top-20 relations and top-2 entities per mention retrieved by our model in Table 6.
Appendix C Baseline Models
@@
  • RetinaQA Faldu et al. (2024) is a two-branch model where one follows a ranking-based approach while the other follows a sketch-filling-based logical form construction method. It then uses a discriminator to determine the final output logical form from the two branches.We note that while we also generate logical form sketches. Such sketches are used for entity mention detection only and is not used to form the final output logical forms directly, i.e., our method is not sketch-filling-based.
@@
  • ChatKBQA Luo et al. (2024a) fine-tunes an open-source LLM to map questions into draft logical forms. The ambiguous KB items in the draft logical forms are replaced with specific KB elements by a separate retrieval module.
@@
  • RoG (Luo et al., 2024b) fine-tunes an LLM to generate relation paths grounded in the KB, given an input question. These generated relation paths are then used to retrieve valid reasoning paths from the KB, enabling faithful reasoning to derive the final answers.
  • ToG (Sun et al., 2024) performs step-by-step relation path selection using an LLM via few-shot in-context learning. At each step, based on newly extended paths, the model prompts the LLM to either answer the question based on the reasoning paths extended so far or continue with another step of path extension.
  • ICSU (Liu et al., 2024) retrieves questions from the training set that are similar to an input question. Pairs of the retrieved questions and the corresponding logical forms are used as examples to prompt an LLM to generate logical forms for the input question.
  • FiDeLiS (Sui et al., 2025) combines semantic similarity metrics with graph-based connectivity to incrementally retrieve and extend reasoning paths. It further proposes Deductive-Verification Beam Search (DVBS), which converts the question into a declarative statement and uses an LLM to verify — by combining this statement with the current tail node of the reasoning path — whether the current node can be directly output as the answer or the path should be further extended.
Appendix D Implementation Details
@@
Following our baselines (Shu et al., 2022; Gu et al., 2023; Faldu et al., 2024), all retrieval and generation models used in our approach are trained separately on the training sets of GrailQA and WebQSP, and then evaluated on their respective validation and test sets. Specifically, for GrailQA, we follow previous works (Shu et al., 2022; Gu et al., 2023; Faldu et al., 2024) and adopt the official data splits described in Section 5.1 to evaluate generalization under non-I.I.D. settings. For WebQSP, we use the same data splits (described in Appendix 5.1) as in previous studies (Shu et al., 2022; Gu et al., 2023; Luo et al., 2024a; Wang and Qin, 2024) to ensure fair comparison.
@@
Appendix E Results on an Additional Evaluation Metric
Model GrailQA WebQSP CWQ
ToG 81.4 82.6 67.6
FiDeLiS - 84.3 71.4
SG-KBQA (ours) 89.0 84.1 82.1
Table 7: Hit results (%) on the validation set of GrailQA and the test set of WebQSP and CWQ.
Following recent studies Sun et al. (2024); Sui et al. (2025), we also compare the Hit results of SG-KBQA with theirs on GrailQA, WebQSP, and CWQ. Although SG-KBQA is not explicitly optimized for producing at least one correct answer, it still shows substantial gains on GrailQA (7.6) and CWQ (10.7) over the baselines. On WebQSP, the Hit of SG-KBQA is only 0.2 lower than the best baselines. These results further confirm the effectiveness of SG-KBQA in enhancing generalizability on non-I.I.D. scenarios and accuracy on complex questions.
Appendix F Parameter Study
We conduct a parameter study to investigate the impact of the choice of values for our system parameters. When the value of a parameter is varied, default values as mentioned in Appendix D are used for the other parameters.
@@ Appendix F Parameter Study
As $k_{L}$ increases, the recall of candidate entity retrieval grows, which is expected. The growth diminishes gradually. This is because a small number of questions contain complex entity mentions that are difficult to handle (see error analysis in Appendix I). As $k_{L}$ increases, the precision of the retrieval also reduces, which brings noise into the entity retrieval results and additional computational costs. To strike a balance, we set $k_{L}=3$ for GrailQA and $k_{L}=4$ for WebQSP. We also observe that the recall on WebQSP is lower than that on GrailQA. This is because WebQSP has a smaller training set to learn from.
@@
SpanMD (span classification): length (✗)
Ours (extracting from generated logical form sketches):
@@
Figure 8: Case study of logical form generation by SG-KBQA and two representative competitors TIARA and PANGU on the GrailQA validation set. Green arrows and boxes highlight schema-level interactions (overlapping classes) among KB elements.
@@
Appendix G Model Running Time
@@
Appendix H Case Study
To further show SG-KBQA’s generalizability to non-I.I.D. KBQA applications, we include a case study from the GrailQA validation set as shown in Table 8 and Figure 8.
Entity Mention Detection
@@ Appendix H Case Study
By retrieving the unseen relation measurement_unit.measurement_system.length_units from the question through our relation retrieval module, the model is provided with schema-level information not encountered during training. This enables it to accurately identify the correct entity mention, atomic units, from surrounding relation tokens—even when both the mention and relation are novel. Moreover, removing schema guidance (i.e., using a Seq2Seq model without SER) results in the same incorrect entity detection as SpanMD, highlighting the importance of schema-guidance in improving entity retrieval in non-I.I.D. scenarios.
@@
Figure 8 shows a logical form generation example. Here, SG-KBQA and TIARA (a representative generation-based model) and PANGU (a representative ranking-based SOTA model) have both retrieved the same sets of relations and entities in the retrieval stage which include false positives and unseen ground-truth relation. Meanwhile, the three models also share the same retrieved entity m.042zlv3. Despite accessing the correct entity and relevant relations, both TIARA and PANGU fail to generate the correct logical form, instead producing a logical form that consists of a relation composition seen at training. In contrast, SG-KBQA successfully generates the correct logical form by leveraging schema context (i.e. the entity’s class as well as the domain and range classes of the retrieved relations which overlap). This enables SG-KBQA to reason about the connectivity between the unseen ground-truth relation and the seen elements, demonstrating its superiority in generalizing to novel compositions of KB elements.
Appendix I Error Analysis